1. Solve the following system of equations using Gauss-Jordan elimination. 3x - 2y +4z=3 2x +2y-2z=4 x+4y- &z=1
Answer #1

Ansi- 30 - 23 + 47 - 3 2 +2% -27 = 4 + 4 -8 - | Now using gauss Jordan ellimination 2. 40 M 웅 R 3 엉 - 2 4. - 8 사 T Ra → 3R2-2Thes is Jordan ellimination form so x - y + y2 = 1 y - 142 = 6 10 Z Now 20 y = 6 + 14 10 10 2 = 1 + 3x2 - 4 11 3+4-4 3 3444